OpenStudy (anonymous):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2, and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold. 1. Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work.

Directrix (directrix):

Slope-intercept form looks like this --> y = m*x + b where m is the slope of the line and b is the y-intercept.

Directrix (directrix):

2x + 3y = 1,200 Subtract 2x from both sides. 2x + 3y = 1,200 -2x -2x ----------------- 3y = -2x + 1200

Directrix (directrix):

Multiply both sides by (1/3) which is the same operation as dividing by 3.

OpenStudy (anonymous):

thank you

Directrix (directrix):

We're not finished.

Directrix (directrix):

y = -2x + 1200 y = (-2/3)x + 400 @autinregister Name the slope and the y-intercept, okay?

OpenStudy (anonymous):

the slope would be (-2/3) and the y-intercept would be 400

Directrix (directrix):

Yes. We say the y-intercept is 400. Its ordered pair would be (0,400).

OpenStudy (anonymous):

Thank you